OpenStudy (diana):

How do you solve cos(2 tan(-4/3))?

OpenStudy (diana):

it's \[\cos (2\tan^{-1} (-4/3))\]

OpenStudy (cornzyblack):

you first solve for 2tan(-4/3) and call it x then you solve for cos(x)

OpenStudy (diana):

but how would you solve that without using a calculator?

OpenStudy (anonymous):

Factor out tan. Remember that tan = sin/cos. If you know the values of sin and cos around a unit circle, you can work the answer out without a calculator. E.g, sin pi/2 = 1, this is because at pi/2 (radian measure for 90 degrees), sin (or y) is 1 UNIT "north" (remember it's a unit circle) of the origin.

OpenStudy (anonymous):

So with this particular question, you'd factor tan (-4/3), which would give you (sin (-4/3)/cos (-4/3)), or (sin -240/cos -240)

OpenStudy (anonymous):

Did that help?

OpenStudy (diana):

wait let me try

OpenStudy (diana):

wait how did you get -240?

OpenStudy (anonymous):

To convert a radian measure into a degree measure, multiply the radian fraction by 180. So what I did was: -4/3 * 180 = -240

OpenStudy (diana):

so what would the answer be?